(2y^2-4x+5)dx+(4-2y+4xy)dy=0
For exactness, and solve it if it is exact.
1
Expert's answer
2020-10-26T19:59:38-0400

Given

"(2y^2-4x+5)dx+(4-2y+4xy)dy=0"

Let "M(x,y)=2y^2-4x+5" and "N(x,y)=4-2y+4xy"

Thus, differential equation becomes

"Mdx+Ndy=0"

Now, to check exactness we have to check

"\partial_yM=\partial_xN"

or not.

In this case ,


"\partial_yM=4y=\partial_xN"

Thus,

"(2y^2-4x+5)dx+(4-2y+4xy)dy=0"

is exact differential equation.

Thus,


"(2y^2-4x+5)dx+(4-2y+4xy)dy=0\\
\implies (-4x+5)dx+(4-2y)dy+d(2xy^2)=0\\
\implies \int(-4x+5)dx+\int(4-2y)dy+\int d(2xy^2)=C\\
\implies -2x^2+5x+4y-y^2+2xy^2=C"